• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Markowitz

    Log entry time 08:11:10 on June 04, 2005

    Entry number 144945

    keyword=owl shift summary

    Beam tuning continued all shift, with accelerator trying to minimize the RF instability and the energy spread. Arne was seeing correlations between beam current and energy spread which he was unable to minimize.

    At 07:00 the accelerator had setup Halls A and B, and asked us to look at the beam quality. Energy spread was good (between 1 and 2e-5, but it was fluctuating). Spot size on arc scanners 5 and 6 was about 60 microns in horizontal direction but over 100 microns in the vertical. Doug Higinbotham is in the MCC helping tune the spot size.

    A copy of this log entry has been emailed to: lerose@jlab.org, garibald@jlab.org